OpenStudy (anonymous):

Write the equation of a hyperbola with vertices at (-3, 0) and (3, 0) and co-vertices (0, 5) and (0, -5).

OpenStudy (anonymous):

x^2/9-y^2/25=1 x^2/3-y^2/5=1 y^2/25-x^2/9=1 y^2/9-x^2/25=1

OpenStudy (anonymous):

@Hero

hero (hero):

@purplemath

OpenStudy (anonymous):

@purplemath ?

OpenStudy (anonymous):

does that exist?

hero (hero):

Yes, it does actually

OpenStudy (anonymous):

ohk

hero (hero):

^correct link

OpenStudy (anonymous):

cool..i just need the answer. but thanks

OpenStudy (anonymous):

@helder_edwin hello?

OpenStudy (anonymous):

@hartnn